A solution containing 30% alcohol is to be mixed with a solution containing 50%
alcohol to make 200 L of a solution containing 42% alcohol. How much of the 50%
solution should be used?

Answer by stanbon(57375)

A solution containing 30% alcohol is to be mixed with a solution containing 50%
alcohol to make 200 L of a solution containing 42% alcohol. How much of the 50%
solution should be used?
----------------------
Equation:
alcohol + alcohol = alcohol
0.30x + 0.50(200-x) = 0.42*200
Multiply thru by 100 to get:
30x + 50*200 - 50x = 42*200
-20x = -8*200
x = 80 liters (amount of 30% solution needed in the mixture)
==================================================================
Cheers,
Stan H.